OpenStudy (anonymous):

4-3x= (blank) x + (blank)-2x

OpenStudy (blurbendy):

4-3x = -x + 4 -2x 4 - 3x = -3x + 4 4 -3x = 4 - 3x

OpenStudy (blurbendy):

make sense?

OpenStudy (anonymous):

you are a life saver!! thank you!!

OpenStudy (blurbendy):

anytime!

OpenStudy (anonymous):

How did u know to choose those numbers

OpenStudy (blurbendy):

the left side had -3x, but on the right side we only had something x - 2x. How do we get -3x on the right side? when that something has to be -1, because -1x -2x = -3x so, all that was missing was the 4, so there was one more blank left, it must be 4

OpenStudy (anonymous):

wow..thank you..

OpenStudy (blurbendy):

of course